IF Y VARIES DIRECTLY WITH X, AND IF Y=30 WHEN X=6, FIND Y WHEN X IS 11
Answer by checkley71(8403)

30/6=Y/11 NOW CROSS MULTIPLY
6Y=30*11
6Y=330
Y=330/6
Y=55 IS THE ANSWER FOR Y WHEN X=11.
PROOF:
30/6=55/11
5=5